# assumption[theorem]Assumption

### v1

@@ 7.1 Constant curvature spaces
which yields the stated scaling of coefficients; the vanishing of odd terms persists on the edge side since the manifold has no boundary.
@@
7.2 Multiplicity realization via generalized Sturm–Liouville (Kreĭn strings)
@@
By Kreĭn’s correspondence for strings, every Stieltjes function arises as the Weyl–Titchmarsh function of a (finite-length, up to reparametrization) Kreĭn string, and the associated spectral measure is exactly $\mu^{+}$; Kreĭn (1952) and the modern treatment in [13, Thm. 4.1]. Applying the affine change $y=\pi-C$ transports $\mu^{+}$ to $\mu_{C}$.
@@
The Kreĭn string realizing $\mu^{+}$ in Theorem 7.5 is unique up to string equivalence (reparametrization of the spatial variable); [13, Thm. 4.1].
Proof 7.8.
Let $m(z)$ denote the Weyl–Titchmarsh function of a (finite-length) Kreĭn string. For a finite string, $m$ is a Stieltjes function and admits an integral representation whose spectral measure is precisely $\mu^{+}$. By [13, Thm. 4.1], the map sending a Kreĭn string (up to equivalence, i.e., reparametrization of the spatial variable) to its Weyl–Titchmarsh function is bijective; in particular, $m(z)$ uniquely determines, and is uniquely determined by, the string up to equivalence. Since $\mu^{+}$ uniquely specifies $m(z)$ (and conversely) via the Stieltjes representation, the realizing Kreĭn string is unique up to equivalence.
@@
Proof 7.10.
The integrability $\int(1+t)^{-1}\,d\mu^{+}(t)<\infty$ holds by the displayed summability, which ensures a finite-length string in the sense of [13, Thm. 4.1]. The rest follows from Theorem 7.5 and the affine change $y=\pi-C$.
@@
Proof 7.12.
This is the classical one-dimensional Weyl law; Titchmarsh [16, Ch. XIII] or Levitan–Sargsjan [12]. Under the stated regularity/positivity, the spectrum is discrete and
@@ 7.3 Spectral clustering remarks
Finally, $\ell(C)\sim\sqrt{\Lambda}\sim\sqrt{|C|/\epsilon}$, hence $m_{\ell(C)}\sim\frac{2}{(d-1)!}\big(\frac{|C|}{\epsilon}\big)^{\frac{d-1}{2}}$, which yields the stated asymptotics.
@@
Proof 7.17.
Under the stated randomization, the event $\{C_{\ell}\in[C-\delta,C]\}$ is equivalent to $\{U\in[0,\delta]\}$ inside a fundamental cell of length $C_{\ell}-C_{\ell+1}$. Since $U$ is uniform, the probability equals $\delta/(C_{\ell}-C_{\ell+1})$. The identity $C_{\ell}-C_{\ell+1}=\epsilon(2\ell+d)$ follows from $\lambda_{\ell+1}-\lambda_{\ell}=2\ell+d$. The large-$\ell$ decay is immediate.
@@ 8.1 The round sphere $S^{d}$
Proof 8.2.
Use $\omega_{d}=\pi^{d/2}/\Gamma(\tfrac{d}{2}+1)$ and $\mathrm{Vol}(S^{d})=2\pi^{(d+1)/2}/\Gamma(\tfrac{d+1}{2})$, then substitute into $\gamma_{d}=\frac{\omega_{d}}{(2\pi)^{d}}\mathrm{Vol}(S^{d})$. For $d=3$ this gives $\gamma_{3}=1/3$. Insert into Theorem 3.1.
@@ 8.2 The flat torus $T^{d}$
Proof 8.6.
Use the standard Weyl law for flat tori [2] and Theorem 3.1.
@@ 8.6 Manifolds with boundary: the ball
$N_{\mu_{C}}(C)\ \sim\ \tfrac{1}{3}\,\epsilon^{-3/2}(\pi-C)^{3/2}-\tfrac{1}{16}\,\epsilon^{-1}(\pi-C)+o(\pi-C).$ (8.11)
@@ 9 Conclusion
The analysis establishes that the edge-variable encoding $C=\pi-\epsilon\lambda$ reproduces Weyl asymptotics with coefficient $\gamma_{d}$ and bulk density exponent (d2)/2(d-2)/2, that this scaling is stable under admissible perturbations, and that for constant curvature spaces the heat coefficients transfer exactly to the edge-side expansion. Multiplicity patterns, although not realizable by smooth Sturm–Liouville operators in dimensions $d>1$, admit a complete representation by Kreĭn strings unique up to equivalence; clustering on spheres averages out in the bulk and does not alter the limiting law. Explicit evaluations of $\gamma_{d}$ for spheres, tori, hyperbolic surfaces, as well as deformed and factored geometries (Berger spheres, lens spaces) and bounded domains, confirm consistency with the reconstruction formula. Open problems include the full extension to general manifolds with boundary, operators beyond the Laplacian, and finer inverse spectral questions.

### v2

@@ 6.1 Constant curvature spaces
For manifolds of constant sectional curvature the Seeley–DeWitt coefficients [10, 11] are universal polynomials in $K$ and $V$, and the heat-transfer identity $H_{\mathrm{edge}}(s)=\Theta_{\Delta}(\epsilon s)$ yields an explicit scaling law for every coefficient.
@@
which yields the stated scaling of coefficients; the vanishing of odd terms persists on the edge side since the manifold has no boundary. ∎
@@
6.2 Kreĭn string realization of the spectral measure
The affine encoding pushes the Riemannian spectrum to a discrete measure $\mu_{C}$ on $(-\infty,\pi)$ with prescribed multiplicities. We prove that this measure is always realizable by a Kreĭn string [12] (a generalized Sturm–Liouville model with measure coefficients), uniquely up to string equivalence [14], and that no classical smooth Sturm–Liouville operator can produce the same spectral growth for $d>1$ [17, 13].
@@ 6.2 Kreĭn string realization of the spectral measure
By Kreĭn’s correspondence for strings, every Stieltjes function arises as the Weyl–Titchmarsh function of a (finite-length, up to reparametrization) Kreĭn string, and the associated spectral measure is exactly $\mu^{+}$; [12] and the modern treatment in [14, Thm. 4.1]. Applying the affine change $y=\pi-C$ transports $\mu^{+}$ to $\mu_{C}$. ∎
@@
The Kreĭn string realizing $\mu^{+}$ in Theorem 6.5 is unique up to string equivalence (reparametrization of the spatial variable); [14, Thm. 4.1].
Proof 6.8.
Let $m(z)$ denote the Weyl–Titchmarsh function of a (finite-length) Kreĭn string. For a finite string, $m$ is a Stieltjes function and admits an integral representation whose spectral measure is precisely $\mu^{+}$. By [14, Thm. 4.1], the map sending a Kreĭn string (up to equivalence, i.e., reparametrization of the spatial variable) to its Weyl–Titchmarsh function is bijective; in particular, $m(z)$ uniquely determines, and is uniquely determined by, the string up to equivalence. Since $\mu^{+}$ uniquely specifies $m(z)$ (and conversely) via the Stieltjes representation, the realizing Kreĭn string is unique up to equivalence. ∎
@@
Proof 6.10.
The integrability $\int(1+t)^{-1}\,d\mu^{+}(t)<\infty$ holds by the displayed summability, which ensures a finite-length string in the sense of [14, Thm. 4.1]. The rest follows from Theorem 6.5 and the affine change $y=\pi-C$. ∎
@@
Proof 6.12.
This is the classical one-dimensional Weyl law; Titchmarsh [17, Ch. XIII] or Levitan–Sargsjan [13]. Under the stated regularity/positivity, the spectrum is discrete and
@@ 6.3 Spectral clustering remarks
Finally, $\ell(C)\sim\sqrt{\Lambda}\sim\sqrt{|C|/\epsilon}$, hence $m_{\ell(C)}\sim\frac{2}{(d-1)!}\big(\frac{|C|}{\epsilon}\big)^{\frac{d-1}{2}}$, which yields the stated asymptotics. ∎
@@
Proof 6.17.
Under the stated randomization, the event $\{C_{\ell}\in[C-\delta,C]\}$ is equivalent to $\{U\in[0,\delta]\}$ inside a fundamental cell of length $C_{\ell}-C_{\ell+1}$. Since $U$ is uniform, the probability equals $\delta/(C_{\ell}-C_{\ell+1})$. The identity $C_{\ell}-C_{\ell+1}=\epsilon(2\ell+d)$ follows from $\lambda_{\ell+1}-\lambda_{\ell}=2\ell+d$. The large-$\ell$ decay is immediate. ∎
@@ 7 Examples
We compute the edge-variable Weyl law explicitly for the principal model geometries, illustrating the dependence of $\gamma_{d}$ on volume and dimension, the universality of the bulk density exponent (d2)/2(d-2)/2, and the robustness of the asymptotics under metric deformations and topological quotients.
@@ 7.1 The round sphere $S^{d}$
The unit round sphere provides the simplest closed example with known spectrum and explicit Weyl constant; the case $d=3$ gives the numerical sanity check $\gamma_{3}=1/3$.
@@
Proof 7.2.
Use $\omega_{d}=\pi^{d/2}/\Gamma(\tfrac{d}{2}+1)$ and $\mathrm{Vol}(S^{d})=2\pi^{(d+1)/2}/\Gamma(\tfrac{d+1}{2})$, then substitute into $\gamma_{d}=\frac{\omega_{d}}{(2\pi)^{d}}\mathrm{Vol}(S^{d})$. For $d=3$ this gives $\gamma_{3}=1/3$. Insert into Theorem 3.1. ∎
@@ 7.2 The flat torus $T^{d}$
For flat tori the Weyl constant depends only on the volume; the bulk exponent (d2)/2(d-2)/2 is unchanged, and the density is asymptotically flat for $d=2$.
@@
Proof 7.6.
Use the standard Weyl law for flat tori [2] and Theorem 3.1. ∎
@@ 7.4 Berger spheres
Berger spheres are metric deformations of $S^{3}$ that alter the low-lying spectrum while leaving the volume unchanged. The following corollary shows that the bulk asymptotics are insensitive to the deformation parameter $k$, illustrating the universality of the rigidity framework.
@@ 7.5 Lens spaces
Lens spaces $L(p,q)=S^{3}/\mathbb{Z}_{p}$ are topological quotients with volume $\mathrm{Vol}(S^{3})/p$. Although lens spaces with different $q$ may be non-isometric but isospectral, the edge-variable bulk asymptotics detect only the volume, confirming that affine encoding preserves Weyl invariants while ignoring finer spectral features.
@@ 7.6 Manifolds with boundary: the ball
Relaxing Assumption 2.1 to allow smooth boundary, we illustrate the two-term bulk expansion of Proposition 3.10 on the unit ball $B^{3}$ with Dirichlet boundary conditions. The boundary term, proportional to the surface area, survives the encoding and appears as the linear subleading correction.
@@
$N_{\mu_{C}}(C)\ \sim\ \tfrac{2}{9\pi}\,\epsilon^{-3/2}(\pi-C)^{3/2}-\tfrac{1}{4}\,\epsilon^{-1}(\pi-C)+o(\pi-C).$ (7.11)
@@ 8 Conclusion
The central result of this paper is a rigidity theorem for spectral encodings: among all polynomial-type reparametrizations $C=\pi-\epsilon\lambda^{k}L(\lambda)$ with $L\in\mathrm{RV}_{0}$, only the affine case $k=1$ preserves the geometric Weyl density exponent (d2)/2(d-2)/2; this uniqueness extends to the full O-regularly varying class, where the bulk power-law forces $\phi\in\mathrm{RV}_{1}$. As a structural consequence of the uniquely correct affine encoding, the edge-variable counting function satisfies $N_{\mu_{C}}(C)\sim\gamma_{d}\,\epsilon^{-d/2}(\pi-C)^{d/2}$ in the bulk, enabling reconstruction of dimension $d$ and Weyl constant $\gamma_{d}$; this transfer is stable under admissible perturbations $\delta=o(\lambda)$ with explicit slowly-varying error rates. The full discrete spectral measure with multiplicities, not realizable by any smooth Sturm–Liouville operator for $d>1$, is completely and uniquely represented by a Kreĭn string; clustering on spheres averages out in the bulk without affecting the limiting exponent. Conceptually, this is organized by asymptotic spectral equivalence classes: encodings with $\phi\in\mathrm{RV}_{k}$ act as dimension-scaling morphisms $d_{\mathrm{as}}\mapsto d_{\mathrm{as}}/k$, and the dimension-preserving endomorphisms are exactly those with $\phi\in\mathrm{RV}_{1}$. Exact heat-trace and zeta-function correspondences confirm coefficient alignment for constant-curvature spaces, and explicit evaluations across spheres, tori, hyperbolic surfaces, Berger spheres, lens spaces, and bounded domains are consistent with the rigidity framework. Open problems include the classification of encodings beyond the O-RV class, extension to manifolds with boundary and to operators beyond the Laplacian, and finer inverse spectral questions at the level of heat invariants.